# The left-to-right minima basis of the group algebra of the symmetric group (updated version)

### v1

@@ 5 Proof of the main theorem
The only remaining addend is the one for $u=w$; this addend equals $\lambda_{w}$ (since we have $\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\operatorname*{cLRM}\nolimits^{\prime}(w)}\,w\right)=\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\beta}w\right)=1$ by Lemma 6 (a)). Thus, the whole sum simplifies to $\lambda_{w}$. Consequently, (18) rewrites as $0=\lambda_{w}$, thus $\lambda_{w}=0$, as desired. This completes the induction step, hence establishing our claim. ∎
@@ 6 Questions
    On the basis $\bigl\{\mathbf{B}_{\operatorname{LRM}^{\prime}(w)}\,w\;\big|\;w\in S_{n}\bigr\}$ from Corollary 2, left multiplication by any element of $\Sigma_{n}$ acts as a triangular matrix. However, right multiplication generally does not. Is there a basis that represents both left and right multiplication by triangular matrices? Theoretical reasoning using the structure of $\Sigma_{n}$ shows that the answer is “yes” [^2] Indeed, if $\mathbf{k}$ is a field of characteristic 0, then the $\mathbf{k}$-algebra $\Sigma_{n}$ becomes a direct product of $\mathbf{k}$’s after we quotient it by its Jacobson radical ([Sch05, §5.2]). Thus, the same is true of the algebra $\Sigma_{n}\otimes\Sigma_{n}^{\operatorname{op}}$. Hence, each left $\Sigma_{n}\otimes\Sigma_{n}^{\operatorname{op}}$-module over a field of characteristic 0 has a basis on which $\Sigma_{n}\otimes\Sigma_{n}^{\operatorname{op}}$ acts by triangular matrices. In other words, each $(\Sigma_{n},\Sigma_{n})$-bimodule over a field of characteristic 0 has a basis on which $\Sigma_{n}$ acts by triangular matrices from both left and right. This applies in particular to the $(\Sigma_{n},\Sigma_{n})$-bimodule $\mathcal{A}$. Using the structure of finitely generated modules over a PID, this can be extended to $\mathbf{k}=\mathbb{Z}$, and thus to all $\mathbf{k}$., but does not suggest a concrete basis. [^3] NB: The analogue of Lemma 1 for the left ideals $\mathcal{L}_{\alpha}:=\mathcal{A}\mathbf{B}_{\alpha}$ instead of the right ideals $\mathcal{R}_{\alpha}$ is false.
@@
    A descent algebra can be defined for any finite Coxeter group [Sol76]. It is desirable to know if an analogue of LRMs, and a basis such as the one in Corollary 2, still exists in this generality. Note that [ChenGongGuo13, §3] defines an analogue of LRMs in type $B$.

### v2

@@ 5 Proof of the main theorem
The only remaining addend is the one for $u=w$; this addend equals $\lambda_{w}$ (since we have $\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\operatorname{cLRM}^{\prime}(w)}\,w\right)=\left[\underline{w}\right]\left(\mathbf{V}_{\beta}\mathbf{B}_{\beta}w\right)=1$ by Lemma 7 (a)). Thus, the whole sum simplifies to $\lambda_{w}$. Consequently, (19) rewrites as $0=\lambda_{w}$, thus $\lambda_{w}=0$, as desired. This completes the induction step, hence establishing our claim. ∎
6 Eigenvalues
We shall now elaborate upon and prove Corollary 3. We begin with notations for left and right multiplications on $\mathcal{A}$:
Definition 7.
Let $\mathbf{a}\in\mathcal{A}$. Then, we let $L\left(\mathbf{a}\right):\mathcal{A}\rightarrow\mathcal{A}$ and $R\left(\mathbf{a}\right):\mathcal{A}\rightarrow\mathcal{A}$ be the maps that send each $\mathbf{x}\in\mathcal{A}$ to $\mathbf{ax}$ and to $\mathbf{xa}$, respectively. These maps $L\left(\mathbf{a}\right)$ and $R\left(\mathbf{a}\right)$ are $\mathbf{k}$-linear, and are known respectively as left and right multiplication by $\mathbf{a}$.
As linear endomorphisms of the free $\mathbf{k}$-module $\mathcal{A}=\mathbf{k}\left[S_{n}\right]$, these maps $L\left(\mathbf{a}\right)$ and $R\left(\mathbf{a}\right)$ can be represented by square matrices with respect to any basis of $\mathcal{A}$, and these matrices have eigenvalues, eigenvectors and characteristic polynomials. In general, their eigenvalues can fail to belong to $\mathbf{k}$; but not so when $\mathbf{a}$ belongs to the descent algebra $\Sigma_{n}$. Namely:
Corollary 4.
[…]
Let $\mathbf{a}=\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}\mathbf{B}_{\alpha}$ (with $\lambda_{\alpha}\in\mathbf{k}$) be an arbitrary element of the descent algebra $\Sigma_{n}$. Then:
  1. (a)
    The map $L\left(\mathbf{a}\right)$ is represented by an upper-triangular matrix with respect to the LRM-basis.
  2. (b)
    The eigenvalues of the map $L\left(\mathbf{a}\right)$ (listed with their algebraic multiplicities) are the numbers $\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}\eta_{\operatorname{cLRM}^{\prime}(w)}\left(\alpha\right)$, where $w$ ranges over $S_{n}$, and where we use the $\eta_{\beta}\left(\alpha\right)$ defined in Theorem 1.
Proof.
(a) The map $L\left(\mathbf{a}\right)$ depends $\mathbf{k}$-linearly on $\mathbf{a}$. Thus, from $\mathbf{a}=\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}\mathbf{B}_{\alpha}$, we obtain $L\left(\mathbf{a}\right)=\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}L\left(\mathbf{B}_{\alpha}\right)$. Hence, it suffices to show that each of the $L\left(\mathbf{B}_{\alpha}\right)$ is represented by an upper-triangular matrix with respect to the LRM-basis. In other words, it suffices to show that for each $\alpha\in\operatorname{Comp}_{n}$ and $v\in S_{n}$, we have
$\left(L\left(\mathbf{B}_{\alpha}\right)\right)\left(\mathbf{B}_{\operatorname{LRM}^{\prime}(v)}\,v\right)$
$=\left(\text{a linear combination of }\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ with }u\trianglelefteq v\right)$ (20)
[…]
$=\eta_{\beta}\left(\alpha\right)\ \mathbf{B}_{\beta}v+\left(\text{a sum of }\mathbf{B}_{\alpha^{\prime}}v\text{ with }\alpha^{\prime}\prec\beta\right).$
In view of
$\left(\text{a sum of }\mathbf{B}_{\alpha^{\prime}}v\text{ with }\alpha^{\prime}\prec\beta\right)$
$\in\sum_{\begin{subarray}{c}\alpha^{\prime}\in\operatorname{Comp}_{n};\\ \alpha^{\prime}\prec\beta\end{subarray}}\mathcal{R}_{\alpha^{\prime}}\qquad\left(\text{since }\mathbf{B}_{\alpha^{\prime}}v\in\mathbf{B}_{\alpha^{\prime}}\mathcal{A}=\mathcal{R}_{\alpha^{\prime}}\right)$
[…]
$=\sum_{\begin{subarray}{c}\alpha^{\prime}\in\operatorname{Comp}_{n};\\ \widetilde{\alpha^{\prime}}\prec_{\pi}\widetilde{\beta}\end{subarray}}\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }\widetilde{\operatorname{cLRM}^{\prime}\left(u\right)}\preceq_{\pi}\widetilde{\alpha^{\prime}}\right\}$
[…]
$=\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }\widetilde{\operatorname{cLRM}^{\prime}\left(u\right)}\prec_{\pi}\widetilde{\beta}\right\}$
[…]
$\subseteq\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }u\vartriangleleft v\right\}$
$\qquad\qquad\left(\text{since }\widetilde{\operatorname{cLRM}^{\prime}\left(u\right)}\prec_{\pi}\widetilde{\beta}=\widetilde{\operatorname*{cLRM}\nolimits^{\prime}\left(v\right)}\text{ entails }u\vartriangleleft v\right),$
this leads to
$\left(L\left(\mathbf{B}_{\alpha}\right)\right)\left(\mathbf{B}_{\beta}\,v\right)\in\eta_{\beta}\left(\alpha\right)\ \mathbf{B}_{\beta}\ v+\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }u\vartriangleleft v\right\}.$
In view of $\mathbf{B}_{\beta}=\mathbf{B}_{\operatorname{LRM}^{\prime}(v)}$ and $\beta=\operatorname{cLRM}^{\prime}\left(v\right)$, this rewrites as
$\left(L\left(\mathbf{B}_{\alpha}\right)\right)\left(\mathbf{B}_{\operatorname{LRM}^{\prime}(v)}\,v\right)$
$\in\eta_{\operatorname{cLRM}^{\prime}\left(v\right)}\left(\alpha\right)\ \mathbf{B}_{\operatorname{LRM}^{\prime}(v)}v+\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }u\vartriangleleft v\right\}$ (21)
$\subseteq\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }u\trianglelefteq v\right\},$
which is precisely (20). Thus, Corollary 4 (a) is proved.
(b) By part (a), the map $L\left(\mathbf{a}\right)$ is represented by an upper-triangular matrix with respect to the LRM-basis. Hence, the eigenvalues of $L\left(\mathbf{a}\right)$ are the diagonal entries of this matrix. But $L\left(\mathbf{a}\right)=\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}L\left(\mathbf{B}_{\alpha}\right)$ (as we saw in the proof of part (a)). Hence, for each $v\in S_{n}$, we have
$\left(L\left(\mathbf{a}\right)\right)\left(\mathbf{B}_{\operatorname{LRM}^{\prime}(v)}\,v\right)$
[…]
$\in\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}\left(\eta_{\operatorname{cLRM}^{\prime}\left(v\right)}\left(\alpha\right)\ \mathbf{B}_{\operatorname{LRM}^{\prime}(v)}v+\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }u\vartriangleleft v\right\}\right)$
$\subseteq\left(\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}\eta_{\operatorname{cLRM}^{\prime}\left(v\right)}\left(\alpha\right)\right)\mathbf{B}_{\operatorname{LRM}^{\prime}(v)}v+\operatorname*{span}\left\{\mathbf{B}_{\operatorname{LRM}^{\prime}(u)}\,u\text{ }\mid\text{ }u\vartriangleleft v\right\}.$
In other words, the diagonal entries of the matrix that represents $L\left(\mathbf{a}\right)$ with respect to the LRM-basis are the scalars $\sum_{\alpha\in\operatorname{Comp}_{n}}\lambda_{\alpha}\eta_{\operatorname{cLRM}^{\prime}(v)}\left(\alpha\right)$ for all $v\in S_{n}$. So these scalars are the eigenvalues of $L\left(\mathbf{a}\right)$. This proves Corollary 4 (b). ∎
Part (b) (but not part (a)) of Corollary 4 holds also for $R\left(\mathbf{a}\right)$ instead of $L\left(\mathbf{a}\right)$. This follows from Corollary 4 (b) via the following fact that has nothing to do with $\Sigma_{n}$:
Proposition 4.
Let $\mathbf{a}\in\mathcal{A}$ be arbitrary. Then:
  1. (a)
    The linear maps $L\left(\mathbf{a}\right)$ and $R\left(\mathbf{a}\right)$ have the same characteristic polynomial, thus the same eigenvalues (with the same algebraic multiplicities).
  2. (b)
    Let $X$ be the matrix representing the map $L\left(\mathbf{a}\right)$ with respect to the basis $\left(w\right)_{w\in S_{n}}$ of $\mathcal{A}$. Let $Y$ be the matrix representing the map $R\left(\mathbf{a}\right)$ with respect to the basis $\left(w^{-1}\right)_{w\in S_{n}}$ of $\mathcal{A}$. Then, $X=Y^{T}$.
Proof.
(b) By linearity, it suffices to prove this when $\mathbf{a}$ is a single permutation $p\in S_{n}$. In this case, the claim boils down to “$pu=v$ if and only if $v^{-1}p=u^{-1}$”. But the latter is a trivial calculation.
(a) This follows from part (b), since mutually transpose matrices have the same characteristic polynomial. ∎
@@ 7 Questions
    On the basis $\bigl\{\mathbf{B}_{\operatorname{LRM}^{\prime}(w)}\,w\;\big|\;w\in S_{n}\bigr\}$ from Corollary 2, left multiplication by any element of $\Sigma_{n}$ acts as a triangular matrix. However, right multiplication generally does not. Is there a basis that represents both left and right multiplication by triangular matrices? Theoretical reasoning using the structure of $\Sigma_{n}$ shows that the answer is “yes” [^2] Indeed, if $\mathbf{k}$ is a field of characteristic 0, then the $\mathbf{k}$-algebra $\Sigma_{n}$ becomes a direct product of $\mathbf{k}$’s after we quotient it by its Jacobson radical ([7, §5.2]). Thus, the same is true of the algebra $\Sigma_{n}\otimes\Sigma_{n}^{\operatorname{op}}$. Hence, each left $\Sigma_{n}\otimes\Sigma_{n}^{\operatorname{op}}$-module over a field of characteristic 0 has a basis on which $\Sigma_{n}\otimes\Sigma_{n}^{\operatorname{op}}$ acts by triangular matrices. In other words, each $(\Sigma_{n},\Sigma_{n})$-bimodule over a field of characteristic 0 has a basis on which $\Sigma_{n}$ acts by triangular matrices from both left and right. This applies in particular to the $(\Sigma_{n},\Sigma_{n})$-bimodule $\mathcal{A}$. Using the structure of finitely generated modules over a PID, this can be extended to $\mathbf{k}=\mathbb{Z}$, and thus to all $\mathbf{k}$., but does not suggest a concrete basis. [^3] NB: The analogue of Lemma 1 for the left ideals $\mathcal{L}_{\alpha}:=\mathcal{A}\mathbf{B}_{\alpha}$ instead of the right ideals $\mathcal{R}_{\alpha}$ is false.
@@
    A descent algebra can be defined for any finite Coxeter group [8]. It is desirable to know if an analogue of LRMs, and a basis such as the one in Corollary 2, still exists in this generality. Note that [2, §3] defines an analogue of LRMs in type $B$.